21Pilots:

a 12.0L container is filled with a gas to a pressure of 2660 torr at 0°C. At what temp. will the pressure inside the container be 3040 torr if volume is held constant?

Nnesha:

Gay Lussac’s law ?

Nnesha:

The law states that pressure is directly proportional to Temperature when the volume is held constant.\[ \large\rm P_1 \alpha T_1 = P_2 \alpha T_2 \] The simplest form of the equation is \[\large\rm \frac{ P_1 }{ T_1 }=\frac{P_2}{T_2}\] (first convert Temp frm Celsius to Kelvin )

21Pilots:

So convert the 0°C to kelvin?

21Pilots:

Which would be 273k.....I think...

Nnesha:

yes

21Pilots:

Then it'd be 2600/273 = 3040/T2?

Nnesha:

looks good

21Pilots:

Okay. I figured it out. Thanks for helping!

Nnesha:

np! :=)) o^_^o